HIGH RESOLUTION SPECTROSCOPY OF THE B̃^{2}A_{1} - X̃^{2}A_{1} TRANSITION OF CaCH_{3} and SrCH_{3}

\underline{P. M. SHERIDAN}, Department of Chemistry, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada; M. J. DICK, Department of Physics, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada; J. G. WANG and P. F. BERNATH, Department of Chemistry, University of Waterloo, 200 University Ave. West, Waterloo, ON, N2L 3G1 Canada.

The B̃^{2}A_{1} - X̃^{2}A_{1} (0_{0}^{0}) bands of CaCH_{3} and SrCH_{3} have been observed at high resolution using laser excitation spectroscopy. The molecules were synthesized in a laser ablation/supersonic expansion spectrometer by the reaction of the ablated metal atoms with a 1% mixture of (CH_{3})_{4}Sn seeded in argon. The spectra for each molecule exhibit a symmetric top structure, with low J lines of multiple K components present. A rotational analysis is currently in progress, and a comparison of the spectroscopic and structural parameters for both molecules will be presented.